# The uniqueness theorem for Kasparov theory

### v1

@@ 4. $KK$-uniqueness
We remind the reader of the standing assumption stated earlier, which still applies to this and the subsequent sections: $G$ denotes an arbitrary second-countable, locally compact group, $\alpha:G\curvearrowright A$ is an action on a separable $\mathrm{C}^{*}$-algebra, and $\beta:G\curvearrowright B$ is a strongly stable action on a $\sigma$-unital $\mathrm{C}^{*}$-algebra.
@@
Let $(\varphi,\mathbbmss{u}):A\to\mathcal{M}(B)$ be a cocycle representation. We define
@@
Let $\alpha:G\curvearrowright A$ be an action on a separable $\mathrm{C}^{*}$-algebra and $\beta:G\curvearrowright B$ a strongly stable action on a $\sigma$-unital $\mathrm{C}^{*}$-algebra. Let $\mathfrak{C}$ be an eligible set of cocycle representations $(A,\alpha)\to(\mathcal{M}(B),\beta)$, and let $(\Phi,\mathbb{U})\in\mathfrak{C}$ be an absorbing element. Given any $U\in\mathcal{U}(\mathscr{D}_{(\Phi,\mathbb{U})})$, we consider the Cuntz pair $\xi(U)\in\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)$ given by $\operatorname{Ad}(U)\circ(\Phi,\mathbb{U})$ and $(\Phi,\mathbb{U})$. Then the assignment $\Xi:K_{1}(\mathscr{D}_{(\Phi,\mathbb{U})})\to\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)/_{\sim_{\mathfrak{C},h}}$ given by $\Xi([U])=\xi(U)$ defines an isomorphism of abelian groups.
@@
Let $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ be two cocycle representations. Suppose that there exists a norm-continuous path of unitaries $U:[0,\infty)\to\mathcal{U}(\mathscr{D}_{(\varphi,\mathbbmss{u})})$ with $U_{0}=\mathbf{1}$ and
@@
This brings us to the main result of the article, which is a generalized dynamical $KK$-uniqueness theorem that improves upon the dynamical stable uniqueness theorem [21, Theorem 6.4]; see [12, Theorems 3.8, 3.10] and [6, Theorem 5.15] for the non-dynamical case. It also entails the uniqueness theorem for nuclear and/or ideal-related $KK$-theory as treated in [18, Theorem 13.9]. The content of this section also represents a somewhat more self-contained approach to answer [6, Question 5.17] in the positive, as well as an assortment of conceivable variants of that question for other versions of $KK$-theory. Restricting its statement to the case of separable $\mathrm{C}^{*}$-algebras, one obtains Introduction.
@@
Let $\alpha:G\curvearrowright A$ be an action on a separable $\mathrm{C}^{*}$-algebra and $\beta:G\curvearrowright B$ a strongly stable action on a $\sigma$-unital $\mathrm{C}^{*}$-algebra. Let $\mathfrak{C}$ be an eligible set of cocycle representations $(A,\alpha)\to(\mathcal{M}(B),\beta)$. Let $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})\in\mathfrak{C}$ be two absorbing elements that form an anchored $(\mathfrak{C};\alpha,\beta)$-Cuntz pair. Then $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})]=0$ in $KK^{G}(\mathfrak{C};\alpha,\beta)$ if and only if $(\varphi,\mathbbmss{u})$ and $(\psi,\mathbbmss{v})$ are strongly asymptotically unitarily equivalent.
@@ 5. A generalized Paschke-type duality
This section is devoted to proving a generalized Paschke-type duality that unifies earlier results of a similar nature. We start with some preparatory observations.
@@
The only thing to show is that this map is surjective. Concretely, we have to show that for every $x\in\mathscr{D}^{w}_{(\varphi,\mathbbmss{u})}$, there exists $y\in\mathscr{D}_{(\varphi,\mathbbmss{u})}$ such that $x-y\in\mathscr{C}_{\varphi}$. Given $x$, we consider the strictly continuous map $\zeta:G\to\mathcal{M}(B)$ given by $\zeta(g)=\beta^{\mathbbmss{u}}_{g}(x)-x$. By definition of $\mathscr{D}^{w}_{(\varphi,\mathbbmss{u})}$, we have $\zeta(G)\varphi(A)\subseteq B$. Furthermore, every map of the form $\zeta(\_)\varphi(a)$ is norm-continuous by section 5. We apply Kasparov’s technical theorem [31, Theorem 1.4] for $B$ in place of $J$ with $G$-action $\beta^{\mathbbmss{u}}$, $\varphi(A)$ in place of $A_{1}$, $\zeta$ in place of $\varphi$, and $A_{2}=\Delta=0$. This yields two positive contractions $M,N\in\mathcal{M}^{\beta^{\mathbbmss{u}}}(B)$ with $M+N=\mathbf{1}$ satisfying $\zeta(G)N\subset B$ and $\varphi(A)M\subset B$. Clearly $M,N\in\mathscr{D}_{(\varphi,\mathbbmss{u})}$. The element $y=xN\in\mathscr{D}_{\varphi}$ then satisfies for all $g\in G$ that
@@
The main idea behind the next lemma and its corollary is analogous to a similar argument in the proof of [45, Lemma 6.4]. We note that the assumption below, which we impose on a cocycle representation, is an analog of (but not identical to) what Thomsen therein refers to as a “saturated” equivariant *-homomorphism.
@@
Below follows our version of the Paschke duality theorem. We note that for $G=\{1\}$, most of the terms simplify and we recover Thomsen’s [44, Theorem 3.2]. His dynamical generalization [45, Theorem 6.2] is very similar, but was phrased in the language of genuinely equivariant representations, whereas the result below is tailored to fit within the framework of cocycle representations.
@@
Let $\mathfrak{C}$ be an eligible set of cocycle representations $(A,\alpha)\to(\mathcal{M}(B),\beta)$ and let $(\Phi,\mathbb{U})\in\mathfrak{C}$ be an absorbing element. The isomorphism $K_{1}(\mathscr{B}_{(\Phi,\mathbb{U})})\cong KK^{G}(\mathfrak{C};\alpha,\beta)$ from section 5 is, strictly speaking, only the first of two possible versions of Paschke duality in our context. It can be viewed as a generalization of [45, Theorem 6.2] in the language of cocycle representations instead of genuinely equivariant ones. The second version of Paschke duality should relate $K_{0}(\mathscr{B}_{(\Phi,\mathbb{U})})$ to a generalized version of the group $KK^{G}_{1}(\alpha,\beta)$ and/or $\operatorname{Ext}^{G}(\alpha,\beta)$ built from cocycle representations or Busby maps related to $\mathfrak{C}$. A rigorous investigation of how such generalized equivariant Ext- or $KK_{1}$-groups would actually look like is beyond the scope of this article, but the apparent template for this kind of second Paschke duality would be [45, Theorems 6.3, 6.5]; see also [44, Section 3]. For $\mathfrak{C}$ being the set of all cocycle representations, the related constructions are available in the literature, and such a result could in this case perhaps be achieved with some extra work if one appeals to [43, Lemma 4.4]. Since that result and its proof appeal to various technical results in the Kasparov picture of $KK^{G}_{1}$ (including the general form of the Kasparov product and its properties from [30]), it is not immediately clear how to generalize the argument to the present context in a self-contained manner.
@@ Acknowledgements
The author was supported by the European Research Council under the European Union’s Horizon Europe research and innovation programme (ERC grant AMEN–101124789). For the purpose of open access, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.

### v2

@@ 4. $KK$-uniqueness
Let $\alpha:G\curvearrowright A$ and $\beta:G\curvearrowright B$ be actions on $\mathrm{C}^{*}$-algebras. Let $(\varphi,\mathbbmss{u}):A\to\mathcal{M}(B)$ be a cocycle representation. We define
@@
We remind the reader of the standing assumptions from subsection 1.1, which still applies to this section and the subsequent section.
@@
The next result is the first of two Paschke-type duality results, as discussed in the introduction. A more refined version of a Paschke duality result, which is of interest mainly for $\mathrm{C}^{*}$-dynamics, will be explored in the next section.
@@
Let $\alpha:G\curvearrowright A$ be an action on a separable $\mathrm{C}^{*}$-algebra and $\beta:G\curvearrowright B$ a strongly stable action on a $\sigma$-unital $\mathrm{C}^{*}$-algebra. Let $\mathfrak{C}$ be an eligible set of cocycle representations $(A,\alpha)\to(\mathcal{M}(B),\beta)$, and let $(\Phi,\mathbb{U})\in\mathfrak{C}$ be an absorbing element. Given any $U\in\mathcal{U}(\mathscr{D}_{(\Phi,\mathbb{U})})$, we consider the Cuntz pair $\xi(U)\in\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)$ given by $\operatorname{Ad}(U)\circ(\Phi,\mathbb{U})$ and $(\Phi,\mathbb{U})$. Then the assignment $\Xi:K_{1}(\mathscr{D}_{(\Phi,\mathbb{U})})\to\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)/_{\sim_{\mathfrak{C},h}}$ given by $\Xi([U])=[\xi(U)]$ defines an isomorphism of abelian groups.
@@
Let $\alpha:G\curvearrowright A$ and $\beta:G\curvearrowright B$ be two actions on $\mathrm{C}^{*}$-algebras such that $A$ is separable and $B$ is $\sigma$-unital. Let $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ be two cocycle representations. Suppose that there exists a norm-continuous path of unitaries $U:[0,\infty)\to\mathcal{U}(\mathscr{D}_{(\varphi,\mathbbmss{u})})$ with $U_{0}=\mathbf{1}$ and
@@
This brings us to the main result of the article, which is a generalized dynamical $KK$-uniqueness theorem that improves upon the dynamical stable uniqueness theorem [21, Theorem 6.4]. It also entails the uniqueness theorem for nuclear and/or ideal-related $KK$-theory as treated in [18, Theorem 13.9]. The content of this section also represents a somewhat more self-contained approach to answer [6, Question 5.17] in the positive, as well as an assortment of conceivable variants of that question for other versions of $KK$-theory. Restricting its statement to the case of separable $\mathrm{C}^{*}$-algebras, one obtains Introduction.
@@
Let $G$ be a second-countable, locally compact group. Let $\alpha:G\curvearrowright A$ be an action on a separable $\mathrm{C}^{*}$-algebra and $\beta:G\curvearrowright B$ a strongly stable action on a $\sigma$-unital $\mathrm{C}^{*}$-algebra. Let $\mathfrak{C}$ be an eligible set of cocycle representations $(A,\alpha)\to(\mathcal{M}(B),\beta)$. Let $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})\in\mathfrak{C}$ be two absorbing elements that form an anchored $(\mathfrak{C};\alpha,\beta)$-Cuntz pair. Then $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})]=0$ in $KK^{G}(\mathfrak{C};\alpha,\beta)$ if and only if $(\varphi,\mathbbmss{u})$ and $(\psi,\mathbbmss{v})$ are strongly asymptotically unitarily equivalent.
@@ 4. $\bullet$
Complementary to section 2, we end this section by recording a number of immediate corollaries that follow from the main result in conjunction with the discussion in section 3.
Corollary 4.7.
Let $A$ be a separable $\mathrm{C}^{*}$-algebra and $B$ a $\sigma$-unital stable $\mathrm{C}^{*}$-algebra. Let $\alpha:G\curvearrowright A$ and $\beta:G\curvearrowright B$ be two actions of a second-countable, locally compact group. Let $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ be two absorbing cocycle representations that form an equivariant $(\alpha,\beta)$-Cuntz pair. If $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})]=0$ in $KK^{G}(\alpha,\beta)$, then $(\varphi,\mathbbmss{u})$ and $(\psi,\mathbbmss{v})$ are strongly asymptotically unitarily equivalent.
Corollary 4.8.
Let $A$ be a separable $\mathrm{C}^{*}$-algebra and $B$ a $\sigma$-unital stable $\mathrm{C}^{*}$-algebra. Let $\varphi,\psi:A\to\mathcal{M}(B)$ be two weakly nuclear and nuclearly absorbing *-homomorphisms that form a Cuntz pair. If $\varphi,\psi]=0$ in $KK_{\mathrm{nuc}}(A,B)$, then $\varphi$ and $\psi$ are strongly asymptotically unitarily equivalent.
Corollary 4.9.
Let $X$ be a topological space. Let $A$ be a separable $X$-$\mathrm{C}^{*}$-algebra and $B$ a $\sigma$-unital stable $X$-$\mathrm{C}^{*}$-algebra. Let $\varphi,\psi:A\to\mathcal{M}(B)$ be two weakly $X$-equivariant *-homomorphisms that form a Cuntz pair.
  1. (i)
    Suppose that $\varphi$ and $\psi$ are both absorbing elements in the set of weakly $X$-equivariant *-homomorphisms. If $\varphi,\psi]=0$ in $KK(X;A,B)$, then $\varphi$ and $\psi$ are strongly asymptotically unitarily equivalent.
  2. (ii)
    Suppose that $\varphi$ and $\psi$ are both weakly nuclear and absorbing elements in the set of weakly nuclear and weakly $X$-equivariant *-homomorphisms. If $\varphi,\psi]=0$ in $KK_{\mathrm{nuc}}(X;A,B)$, then $\varphi$ and $\psi$ are strongly asymptotically unitarily equivalent.
@@ 5. A generalized Paschke-type duality
This section is devoted to proving a generalized Paschke-type duality for $\mathrm{C}^{*}$-dynamics that unifies earlier results of a similar nature. We note, however, that the content of this section is vacuous for nonequivariant $KK$-theory, i.e., the case $G=\{1\}$, since it does not add anything beyond the statement of section 4. We start with some preparatory observations.
@@
The only thing to show is that this map is surjective. Concretely, we have to show that for every $x\in\mathscr{D}^{w}_{(\varphi,\mathbbmss{u})}$, there exists $y\in\mathscr{D}_{(\varphi,\mathbbmss{u})}$ such that $x-y\in\mathscr{C}_{\varphi}$. Given $x$, we consider the strictly continuous map $\zeta:G\to\mathcal{M}(B)$ given by $\zeta(g)=\beta^{\mathbbmss{u}}_{g}(x)-x$. By definition of $\mathscr{D}^{w}_{(\varphi,\mathbbmss{u})}$, we have $\zeta(G)\varphi(A)\subseteq B$. Furthermore, every map of the form $\zeta(\_)\varphi(a)$ is norm-continuous by section 5. We apply Kasparov’s technical theorem [30, Theorem 1.4] for $B$ in place of $J$ with $G$-action $\beta^{\mathbbmss{u}}$, $\varphi(A)$ in place of $A_{1}$, $\zeta$ in place of $\varphi$, and $A_{2}=\Delta=0$. This yields two positive contractions $M,N\in\mathcal{M}^{\beta^{\mathbbmss{u}}}(B)$ with $M+N=\mathbf{1}$ satisfying $\zeta(G)N\subset B$ and $\varphi(A)M\subset B$. Clearly $M,N\in\mathscr{D}_{(\varphi,\mathbbmss{u})}$. The element $y=xN\in\mathscr{D}_{\varphi}$ then satisfies for all $g\in G$ that
@@
The main idea behind the next lemma and its corollary is analogous to a similar argument in the proof of [44, Lemma 6.4]. We note that the assumption below, which we impose on a cocycle representation, is an analog of (but not identical to) what Thomsen therein refers to as a “saturated” equivariant *-homomorphism.
@@
Below follows our more refined version of the Paschke duality theorem. We note that for $G=\{1\}$, most of the terms simplify and then it just amounts to the statement of section 4, and we recover Thomsen’s [43, Theorem 3.2]. His dynamical generalization [44, Theorem 6.2] is very similar, but was phrased in the language of genuinely equivariant representations, whereas the result below is tailored to fit within the framework of cocycle representations.
@@
Let $\mathfrak{C}$ be an eligible set of cocycle representations $(A,\alpha)\to(\mathcal{M}(B),\beta)$ and let $(\Phi,\mathbb{U})\in\mathfrak{C}$ be an absorbing element. The isomorphism $K_{1}(\mathscr{B}_{(\Phi,\mathbb{U})})\cong KK^{G}(\mathfrak{C};\alpha,\beta)$ from section 5 is, strictly speaking, only the first of two possible versions of Paschke duality in our context. It can be viewed as a generalization of [44, Theorem 6.2] in the language of cocycle representations instead of genuinely equivariant ones. The second version of Paschke duality should relate $K_{0}(\mathscr{B}_{(\Phi,\mathbb{U})})$ to a generalized version of the group $KK^{G}_{1}(\alpha,\beta)$ and/or $\operatorname{Ext}^{G}(\alpha,\beta)$ built from cocycle representations or Busby maps related to $\mathfrak{C}$. A rigorous investigation of how such generalized equivariant Ext- or $KK_{1}$-groups would actually look like is beyond the scope of this article, but the apparent template for this kind of second Paschke duality would be [44, Theorems 6.3, 6.5]; see also [43, Section 3]. For $\mathfrak{C}$ being the set of all cocycle representations, the related constructions are available in the literature, and such a result could in this case perhaps be achieved with some extra work if one appeals to [42, Lemma 4.4]. Since that result and its proof appeal to various technical results in the Kasparov picture of $KK^{G}_{1}$ (including the general form of the Kasparov product and its properties from [29]), it is not immediately clear how to generalize the argument to the present context in a self-contained manner.
@@ Acknowledgements
The author was supported by the European Research Council under the European Union’s Horizon Europe research and innovation programme (ERC grant AMEN–101124789). For the purpose of open access, the author applies a CC BY public copyright license to any author accepted manuscript version arising from the submission of this work.